If you invested $ 2,000 at 8%, what is the FUTURE VALUE after 3 years?

  1. If you invested $ 2,000 at 8%, what is the FUTURE VALUE after 3 years?
Answer #1

Ans $ 2519.42

If you invested $ 2,000 at 8%, the FUTURE VALUE after 3 years would be $ 2519.42.

FV = Future Value PV = Present Value r= rate of interest n= no of period FV = PV (1+r)^n FV = 2000*(1+8%)^3 FV = 2519.42